Home‎ > ‎Crew 149‎ > ‎

212-CommandersReport

Commander’s Report  
2/12/2015
Paul Bakken

1 – Cross crew support
2 – Comms
3 – Additive manufacturing set up
4 – Geological team EVA
5 – Ballistic launched aerial camera test
6 – Regolith growth study
7 – Lichen study
8 – Food evaluation initiated

Greetings Mission Support!

Life on Mars is starting to feel normal.  Crew woke up early today and immediately set about our daily tasks.

1 – Crew 149 continues to support the projects of other crews.  Ann-Sofie Schreurs cared for Cyprien Verseux’s cyanobacteria soil project.  The little radishes are adorable.  We continue to collect data for Lucie Poulet’s GreenHab activity study.

2 – We investigated a number of avenues to resolve comms.  We searched the Hab unsuccessfully for a substitute router.  We also investigated the possible use of cell signals as an alternate comm channel during EVA 149-3.

3 – The 3D printer was set up in the engineering lock and a test print of a wrench initiated.  We have our eyes on a loose bolt on the loft ladder that may be the first target for the wrench.

4 – Crew members Bakken, Nicoletatos, and Gerardi executed a successful EVA (EVA 149-2) to survey the areas to the SE of the Hab, within 1 km, for deposits of construction grade mineral resources.  The team also conducted a test of the aerial camera.

5 – The personnel of EVA 149-2 also tested the ballistic launch aerial camera.  The results were mostly comical, but somewhat informative.  The padding system works very well to protect the camera upon litho-breaking, but the stabilization system does not work at all.  Immediately after launch, the camera tumbles so rapidly that the video footage is not useable.  The camera teams will explore options for modifying the design to provide better in-flight stabilization.

6 – A number of the hops rhizomes have sprouted new bines!  Overall, the hops seem to be adapting to the regolith growth medium well.  No germination of sorghum in either the experimental or control group yet.  Crew members Nicoletatos and Gerardi completed the renaming and re-labelling of the study plants. (The original labels started to peel off.)

7 – Crew members Schreurs and Naganuma prepared samples of lichens collected during EVA 149-1 for future analysis.  They appear to have collected lichens every color of the rainbow.  Crew has taken to referring to them as our Martian Chia Pets.

8 – We started the food evaluation study today.  Cdr. Bakken ate a few variations of his custom formulated Generic Lipo-enhanced Organic Paste (GLOP) for breakfast and lunch.  Other crew sampled of this unique fare and also filled out evaluation forms.  Version #1 was universally pronounced “dreadful.”  The kindest evaluation opined, “A person could probably live off of this, but it wouldn’t be worth living if they had to.”  Version #2 received higher marks, being somehow both “bland” and chocolatey” at the same time.  Version #3 added a bit of mint extract to Version #2, and received the highest ratings.  Version #4 added orange extract the Version #2, and was disappointingly flat.  Further investigation continues, but it look like chocolate and mint might be the ticket.


Once again, the crew is calling me to supper as I am writing my report, so I will close.  The chicken and beef fried rice looks heavenly.

Respectfully submitted,

Paul Bakken
Commander, MDRS Crew 149